State the slope and graph the line:
1) 4x-7=13
2) 13+2y=5
3) 2x-7y+12=0

Answer by mananth(16949)

1) 4x-7=13
4x=13+7
4x=20
x=5
it ia a line parallel to y axis hence slope is not definable.

2) 13+2y=5
2y=-13+5
2y=-8
y=-4
it is a line parallel to the x axis. so slope is not defined
3) 2x-7y+12=0
7y=2x-12
y=(2x/7)-12/7
slope = 2/7
...
m.ananth@hotmail.ca